# Nonparametric Identification and Estimation of Causal Effects on Latent Outcomes

### 09 Apr 2026 v1

@@ 1 Introduction
Our identification strategy parallels recent bridge-function approaches in the proximal causal inference literature, where auxiliary variables are used to recover information about unobserved quantities (Miao et al., 2018, 2024; Tchetgen et al., 2020). Our contribution is to extend this logic to causal inference with latent outcomes. We propose using measurement bridge functions to align imperfect measurements to a common benchmark, so that downstream causal analysis targets a comparable latent estimand. The bridge function is identified through nonparametric instrumental variables, which constitutes an ill-posed inverse problem and poses challenges for estimation (Newey and Powell, 2003; Darolles et al., 2011; Horowitz, 2011; Chen et al., 2025). The problem becomes more tractable when the target parameter is a linear functional of the nonparametric function (Severini and Tripathi, 2012; Santos, 2011; Zhang et al., 2023). Our estimation approach builds primarily on Bennett et al. (2025).
Our paper also relates to a large literature on the measurement of latent variables. A central insight in this literature is that latent constructs have no intrinsic metric and must be defined through their relationships with observed indicators; they are typically measured with error. Classical work in SEM formalizes this idea by specifying a measurement model that links observed outcomes to an unobserved latent variable, typically through factor loadings and measurement errors (Bagozzi, 1977; Sorbom, 1981; Kano, 2000). Item response theory (IRT) is also widely used in the social sciences (Embretson and Reise, 2013; Zhou, 2019; Lord and Novick, 2008). In contrast to these explicit modeling approaches, when multiple measurements are available, ICW is a dominant method for efficient hypothesis testing (Anderson and Gerbing, 1988). Another commonly used approach is PCA (Hotelling, 1933). We argue that when there is a well-defined latent outcome, our method can yield more informative estimates.
@@ 2 Framework
Consider a sample of $n$ units drawn from a superpopulation. Let $Z_{i}$ denote the binary treatment assignment for unit $i$. The outcome of interest is latent: $\eta_{i}^{1}$ is the latent outcome that would be realized if unit $i$ is assigned to receive the treatment, and $\eta_{i}^{0}$ is the latent outcome that would be realized in the absence of treatment. We assume that the stable unit treatment value assumption (SUTVA) holds (Rubin, 1980). [^3] The SUTVA assumption is a standard requirement for any study of cause and effect; in this context, it implies that potential outcomes remain stable regardless of which subjects receive treatment, ruling out, for example, spillovers between subjects. Formally, (1) If $Z_{i}=Z^{\prime}_{i}$, then $\eta_{i}(Z_{1},Z_{2},...,Z_{i},...,Z_{n})=\eta_{i}(Z_{1},Z_{2},...,Z^{\prime}_{i},...,Z_{n})$; (2) If $Z_{i}=z$, then $\eta_{i}=\eta_{i}(z)$, $\forall i$ and $z\in Z$. These latent variables are not directly observed. Examples include ideology and state capacity in political science, preferences and human capital in economics, mental health or cognitive ability in psychology, and socioeconomic status and social capital in sociology. These constructs are conceptually important and constitute the true outcomes of interest.
@@ 3.2 Measurement Noncomparability Challenge
3.3 Our Solution: Nonparamteric Scaled Index (NSI)
A satisfactory estimator must therefore address two challenge simultaneously: it must render different measurements comparable within a study while remaining robust to model misspecification and preserving information, and it must ensure that causal estimands are comparable across studies.
First, because our object of interest is the ALTE and the multiple measurements are designed to capture the same latent outcome, it is desirable to extract a common latent signal from these measurements. Ignoring the latent structure would result in a loss of useful information. Moreover, because the latent outcome is measurement-dependent, ignoring the latent structure may exacerbate the study noncomparability challenge. However, measurement noncomparability suggests that we require a bridge function that transforms different measurements so that they convey the same information, at least in expectation, before extracting the common latent signal. Formally, we require a measurement bridge function $\varphi_{j}$ for each measurement $Y_{j}$ such that $\mathbb{E}[\varphi_{j}(Y_{k})|\eta]=\mathbb{E}[\varphi_{k}(Y_{k})|\eta]$.
Next, to address the study noncomparability challenge, the dimension-reduction mapping $M$ must preserve the common information shared across different sets of measurements. This is a strong requirement, as it must hold for all possible measurement sets, while the latent outcome is, in principle, defined through multiple measurements. Rather than imposing strong restrictions on the mapping $M$, we propose a design-based approach. Specifically, we require the existence of a common measurement—referred to as a bridge variable—that links different sets of measurements.
@@ 3.3 Our Solution: Nonparamteric Scaled Index (NSI)
Choose Benchmark $Y_{1}$ Estimate Bridge Function $\varphi_{j}(Y_{j})$ Downsteam Causal Analysis
@@
We then use these nonparametrically transformed measurements for downstream causal analysis. For example, we can combine the transformed measurements through a weighted average, $\tilde{Y}=\sum_{j=1}^{J}\omega_{j}\varphi_{j}(Y_{j})$, with weights satisfying $\sum_{j=1}^{J}\omega_{j}=1$. The optimal weights that minimize variance can be used. We then treat $\tilde{Y}$ as the new observed outcome to estimate average treatment effects.
@@
For the proposed method to be meaningful—so that $\tilde{Y}$ captures the latent variable $\eta$—we impose the following centering assumption.
@@
There exists at least one measurement $Y_{j}$ such that is centering in the sense $\mathbb{E}[Y_{j}|\eta]=\eta$.
Without loss of generality, we take the centering variable to be $Y_{1}$. This assumption implies that the measurement provides an unbiased representation of the latent outcome. Many methods have been proposed to obtain unbiased latent variable from biased measurement (Vishnubhatla et al., 2026). If multiple centering measurements exist, the choice of benchmark variable is inconsequential. We emphasize that only one centering measurement is required, and we allow the remaining measurements to have arbitrary (possibly nonlinear) relationships with the latent variable $\eta$. For these non-centering measurements, we use a bridge function $\varphi_{j}$ to transform them so that they are centered in the same way as $Y_{1}$.
@@ 4.1 Existence of the Measurement Bridge Function
From functional analysis, the measurement bridge function $\varphi$ exists if and only if $r$ lies in the range of the operator $K$. In general, this condition is not easy to verify. We therefore provide a sufficient condition that has practical implications.
@@
i. $\mathbb{E}[Y^{2}_{ij}|\eta_{i}]<\infty$. ii. $\iint f(y_{ij}|\eta_{i})f(\eta_{i}|y_{ij})dy_{ij}d\eta_{i}<\infty$ iii. $\sum_{n=1}^{\infty}\frac{1}{\mu_{n}^{2}}|<r,\psi_{n}>|^{2}<\infty$
@@ 4.2 Identification
(2) (Completeness) For any square-integrable function $g$, $\mathbb{E}[g(Y_{i1})|W_{i}]=0$ for almost all $W_{i}$ implies $g(Y_{i1})=0$ almost surely
@@
Once $\varphi_{j}$ is identified, we obtain transformed measurements $\tilde{Y}j=\varphi_{j}(Y{ij})$. In this case, we have multiple outcome variables, so that the ALTE is over-identified. There are several ways to utilize these multiple measurements. For example, we can combine the $J$ transformed measurements to form a single “observed” outcome variable $\tilde{Y}_{i}=\sum_{j=1}^{J}\omega_{j}\tilde{Y}_{j}$, where the weights satisfy $\sum_{j=1}^{J}\omega_{j}=1$. To achieve minimal variance, we propose using inverse-variance weights, $\omega_{j}=\frac{\sigma_{j}^{2}}{\sum_{j=1}^{J}\sigma^{2}_{j}}$, where $\sigma_{j}^{2}$ denotes the variance of the transformed outcome. With the optimally weighted outcome $\tilde{Y}_{i}$, we can then conduct standard causal inference. The average LTE is identified as $\mathbb{E}[\tilde{Y}_{i}\mid Z=1]-\mathbb{E}[\tilde{Y}_{i}\mid Z=0]$. Alternatively, a more convenient approach is to use optimal GMM to estimate the ALTE.
@@ 4.3 Estimation Under Weak Identification
$\Xi_{0}=argmin_{\xi_{j}\in\Phi_{j}}\frac{1}{2}\mathbb{E}[\mathbb{E}[\xi_{j}(Y_{j})|W]^{2}]-\mathbb{E}[s(Z_{i})\phi_{j}(Y_{j})]$
This assumption essentially impose on the smoothness of ALTE $\tau=\mathbb{E}[s(Z_{i})\phi_{j}(Y_{j})]$ with respect to the NPIV conditional expectation operator $K$. Similar to the DML literature (Chernozhukov et al., 2018), the key step is to construct a Neyman orthogonal score for the ALTE. The Neyman orthogonal score is constructed by introducing a debiasing nuisance function $q^{*}_{j}(W)=\mathbb{E}[\xi_{0}(Y_{j})\mid W]$. Then, the score for the ALTE is given by $\psi_{j}=\alpha\phi_{j}(Y_{j})+q_{j}(W)(Y_{1}-\phi_{j}(Y_{j}))-\tau$. There are three parameters in the score that must be estimated. To avoid overfitting bias, all nuisance functions are estimated via cross-fitting. Let the sample be split into $K$ folds $I_{1},\ldots,I_{K}$. For each fold $k$, we estimate the bridge function using a penalized minimax criterion:
@@
$\sqrt{n}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{m}_{ij}-\tau_{0}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{ij}+op(1),\qquad j=2,\ldots,J.$
@@ 4.4 Simulation
In $1000$ Monte Carlo simulations with sample size $n=800$ per study, we compute the difference in ALTE across two studies and conduct a Wald test of equality between the two estimates. The results are reported in Table 1. The average cross-study gap is $0.256$ for PCA and $0.366$ for ICW. PCA rejects the true null hypothesis of equal ALTE in $27.4\%$ of replications, while ICW rejects it in $100\%$ of replications. As expected, because PCA and ICW are measurement-dependent methods, the resulting estimates are also measurement-dependent and therefore not comparable, even when the latent outcome and ALTE are identical across studies.
@@ 5 Application
Table 2: Bennett regression estimates for the Kalla–Broockman application
@@
  • Notes: Entries are coefficient estimates with standard errors in parentheses. $^{*}p<0.10$, $^{**}p<0.05$, and $^{***}p<0.01$. The first three columns report Bennett estimators using series, random-forest, and RKHS bases, respectively, to estimate the bridge component. In the Bennett implementation, the instrument set includes the two treatment indicators and baseline covariate. The last column reports the weighted scaled index (WSI) estimator from the linear model benchmark.
Table 2 reports the results. We use two treatment assignment and baseline covariate as IVs in the first stage to estimate the bridge function. Panel A presents specifications in which the latent outcome is regressed on the two treatment indicators only. Panel B adds the baseline covariate to the second-stage regression. Across all three nonparametric implementations, the estimated effect of the full treatment is positive and statistically distinguishable from zero at conventional levels in most specifications. The point estimates are also quite stable across basis choices, ranging from $0.418$ to $0.444$ without covariate adjustment and from $0.398$ to $0.425$ with covariate adjustment. The abbreviated treatment has a smaller estimated effect and is not statistically distinguishable from zero in either panel. These findings mirror the substantive conclusion in Fu and Green (2025): the full canvassing intervention produces a meaningful shift in the latent immigration-attitude outcome, whereas the abbreviated treatment does not.
@@ 1. $r$ belongs to the orthogonal complement $\mathcal{N}(K^{*})^{\perp}$; and
We observe that $\mathbb{E}[\varphi_{j}(Y_{ij})|\eta_{i}]=\mathbb{E}[Y_{i1}|\eta_{i}]$ can be represented by $K\varphi=r$, where $r(\eta_{i}):=\mathbb{E}[Y_{ij}|\eta_{i}]\in H_{2}$ by regularity assumption i.
@@ B.2 Proof of Proposition 2
Now, let $K:\varphi\mapsto\mathbb{E}[\varphi\mid W_{i}]$ and define $r(W_{i})=\mathbb{E}[Y_{i1}\mid W_{i}]$. We obtain the well known NPIV equation $K\varphi=r$. The function $\varphi$ uniquely identified if and only if completeness condition hold: for any square-integrable function $g$, $\mathbb{E}[g(Y_{i1})|W_{i}]=0$ for almost all $W_{i}$ implies $g(Y_{i1})=0$ almost surely (Newey and Powell, 2003). ∎

### 21 Apr 2026 v2

@@ 1 Introduction
Our identification strategy parallels recent bridge-function approaches in the proximal causal inference literature, where auxiliary variables are used to recover information about unobserved quantities (Miao et al., 2018, 2024; Tchetgen et al., 2020). Our contribution is to extend this logic to causal inference with latent outcomes. We propose using measurement bridge functions to align imperfect measurements to a common benchmark, so that downstream causal analysis targets a comparable latent estimand. The bridge function is identified through nonparametric instrumental variables, which constitutes an ill-posed inverse problem and poses challenges for estimation (Newey and Powell, 2003; Darolles et al., 2011; Horowitz, 2011; Chen et al., 2025). The problem becomes more tractable when the target parameter is a linear functional of the nonparametric function (Chen and Gao, 2026; Severini and Tripathi, 2012; Santos, 2011; Zhang et al., 2023). Blundell et al. (2007) was among the first to use bounded completeness for the identification of NPIV models and to develop efficient estimation methods for finite-dimensional parameters. Our estimation approach builds primarily on Bennett et al. (2025).
Our paper also relates to a large literature on the measurement of latent variables. A central insight in this literature is that latent constructs have no intrinsic metric and must be defined through their relationships with observed indicators; they are typically measured with error. Classical work in SEM formalizes this idea by specifying a measurement model that links observed outcomes to an unobserved latent variable, typically through factor loadings and measurement errors (Bagozzi, 1977; Sorbom, 1981; Kano, 2000). Item response theory (IRT) is also widely used in the social sciences (Embretson and Reise, 2013; Zhou, 2019; Lord and Novick, 2008). In contrast to these explicit modeling approaches, when multiple measurements are available, ICW is a dominant method for efficient hypothesis testing (Anderson and Gerbing, 1988). Another commonly used approach is PCA (Hotelling, 1933). Multiple measurements and factor models are also broadly discussed in the measurement error literature (see the review by Schennach (2022)). We argue that when there is a well-defined latent outcome, our method can yield more informative estimates.
@@ 2 Framework
Consider a sample of $n$ units drawn from a superpopulation. Let $Z_{i}$ denote the binary treatment assignment for unit $i$. The outcome of interest is latent: $\eta_{i}^{1}$ is the latent outcome that would be realized if unit $i$ is assigned to receive the treatment, and $\eta_{i}^{0}$ is the latent outcome that would be realized in the absence of treatment. We assume that the stable unit treatment value assumption (SUTVA) holds (Rubin, 1980). [^3] The SUTVA assumption is a standard requirement for any study of cause and effect; in this context, it implies that potential outcomes remain stable regardless of which subjects receive treatment, ruling out, for example, spillovers between subjects. Formally, (1) $\eta_{i}(z_{i},z_{-i})=\eta_{i}(z_{i},z^{\prime}_{-i})\forall z_{-i},z^{\prime}_{-i}$; (2) If $Z_{i}=z$, then $\eta_{i}=\eta_{i}(z)$, $\forall i$ and $z\in Z$. These latent variables are not directly observed. Examples include ideology and state capacity in political science, preferences and human capital in economics, mental health or cognitive ability in psychology, and socioeconomic status and social capital in sociology. These constructs are conceptually important and constitute the true outcomes of interest.
@@ 3.2 Measurement Noncomparability Challenge
3.3 Our Solution: Nonparametric Scaled Index (NSI)
A satisfactory estimator must therefore address two challenges simultaneously: it must render different measurements comparable within a study while remaining robust to model misspecification and preserving information, and it must ensure that causal estimands are comparable across studies.
First, because our object of interest is the ALTE and the multiple measurements are designed to capture the same latent outcome, it is desirable to extract a common latent signal from these measurements. Ignoring the latent structure would result in a loss of useful information. Moreover, because the latent outcome is measurement-dependent, ignoring the latent structure may exacerbate the study noncomparability challenge. However, measurement noncomparability suggests that we require a bridge function that transforms different measurements so that they convey the same information, at least in expectation, before extracting the common latent signal. Formally, we require a measurement bridge function $\varphi_{j}$ for measurement $Y_{j}$, so that $\mathbb{E}[\varphi_{j}(Y_{j})\mid\eta]=\mathbb{E}[Y_{1}\mid\eta]$. Therefore, conditional on latent outcome $\eta$, all measurements on expectation are equal to $Y_{1}$. $Y_{1}$ here is chosen arbitrarily.
Next, to address the study noncomparability challenge, the dimension-reduction mapping $M$ must preserve the common information shared across different sets of measurements. This is a strong requirement, as it must hold for all possible measurement sets, while the latent outcome is, in principle, defined through multiple measurements. Rather than imposing strong restrictions on the mapping $M$, we propose a design-based approach. Specifically, we require the existence of a common measurement across different studies. Without loss of generality, we denote this as the first measurement, $Y_{1}$. This variable serves as a bridge linking different sets of measurements.
@@ 3.3 Our Solution: Nonparametric Scaled Index (NSI)
Choose Benchmark $Y_{1}$ Estimate Bridge Function $\varphi_{j}(Y_{j})$ Downstream Causal Analysis
@@
Because the latent outcome has no intrinsic scale or unit, we interpret it using the same metric of $Y_{1}$. In other words, we treat $Y_{1}$ and the nonparametrically transformed measurements as outcome variables. We then use them for downstream causal analysis. For example, we can combine the transformed measurements through a weighted average, $\tilde{Y}=\sum_{j=1}^{J}\omega_{j}\varphi_{j}(Y_{j})$, with weights satisfying $\sum_{j=1}^{J}\omega_{j}=1$. The optimal weights that minimize variance can be used. We then treat $\tilde{Y}$ as the new observed outcome to estimate average treatment effects.
@@
For the proposed method to be meaningful—so that $Y_{1}$ captures the latent variable $\eta$—we can impose the following centering assumption.
@@
Measurement $Y_{1}$ is centering in the sense $\mathbb{E}[Y_{1}|\eta]=\eta$.
This assumption implies that the measurement provides an unbiased representation of the latent outcome. Many methods have been proposed to obtain unbiased latent variable from biased measurement (Vishnubhatla et al., 2026). If multiple centering measurements exist, the choice of benchmark variable is inconsequential. We emphasize that only one centering measurement is required, and we allow the remaining measurements to have arbitrary (possibly nonlinear) relationships with the latent variable $\eta$. For these non-centering measurements, we use a bridge function $\varphi_{j}$ to transform them so that they are centered in the same way as $Y_{1}$.
@@ 4.1 Existence of the Measurement Bridge Function
From functional analysis, the measurement bridge function $\varphi$ exists if and only if $r$ lies in the range of the operator $K$. In general, this condition is not easy to verify. The common sufficient condition that has practical implications is the completeness condition.
@@
i. $\mathbb{E}[Y^{2}_{i1}|\eta_{i}]<\infty$. ii. $\iint f(y_{ij}|\eta_{i})f(\eta_{i}|y_{ij})dy_{ij}d\eta_{i}<\infty$ iii. $\sum_{n=1}^{\infty}\frac{1}{\mu_{n}^{2}}|<r,\psi_{n}>|^{2}<\infty$
@@ 4.2 Identification
(2) (Completeness) For any square-integrable function $g$ and $j\in\{1,...J\}$, $\mathbb{E}[g(Y_{ij})|W_{i}]=0$ for almost all $W_{i}$ implies $g(Y_{ij})=0$ almost surely
@@
Once $\varphi_{j}$ is identified, we obtain transformed measurements $\tilde{Y}_{j}=\varphi_{j}(Y_{ij})$. In this case, we have multiple outcome variables, so that the ALTE is over-identified. There are several ways to utilize these multiple measurements. For example, we can combine the $J$ transformed measurements to form a single “observed” outcome variable $\tilde{Y}_{i}=\sum_{j=1}^{J}\omega_{j}\tilde{Y}_{j}$, where the weights satisfy $\sum_{j=1}^{J}\omega_{j}=1$. To achieve minimal variance, we propose using inverse-variance weights, $\omega_{j}=\frac{\Sigma^{-1}I_{J}}{I^{{}^{\prime}}_{J}\Sigma^{-1}I_{J}}$, where $\Sigma$ denotes the variance-covariance matrix of the transformed outcome. With the optimally weighted outcome $\tilde{Y}_{i}$, we can then conduct standard causal inference. The average LTE is identified as $\mathbb{E}[\tilde{Y}_{i}\mid Z=1]-\mathbb{E}[\tilde{Y}_{i}\mid Z=0]$. Alternatively, a more convenient approach is to use optimal GMM to estimate the ALTE.
@@ 4.3 Estimation Under Weak Identification
$\Xi_{0}=argmin_{\xi_{j}\in\Phi_{j}}\frac{1}{2}\mathbb{E}[\mathbb{E}[\xi_{j}(Y_{j})|W]^{2}]-\mathbb{E}[s(Z_{i})\xi_{j}(Y_{j})]$
This assumption essentially impose on the smoothness of ALTE $\tau=\mathbb{E}[s(Z_{i})\phi_{j}(Y_{j})]$ with respect to the NPIV conditional expectation operator $K$. Similar to the DML literature (Chernozhukov et al., 2018), the key step is to construct a Neyman orthogonal score for the ALTE. The Neyman orthogonal score is constructed by introducing a debiasing nuisance function $q^{*}_{j}(W)=\mathbb{E}[\xi_{0}(Y_{j})\mid W]$. Then, the score for the ALTE is given by $\psi_{j}=s(Z_{i})\phi_{j}(Y_{j})+q_{j}(W)(Y_{1}-\phi_{j}(Y_{j}))-\tau$. There are three parameters in the score that must be estimated. To avoid overfitting bias, all nuisance functions are estimated via cross-fitting. Let the sample be split into $K$ folds $I_{1},\ldots,I_{K}$. For each fold $k$, we estimate the bridge function using a penalized minimax criterion:
@@
$\sqrt{n}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{m}_{ij}-\tau_{0}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{ij}+o_{p}(1),\qquad j=2,\ldots,J.$
@@ 4.4 Simulation
In $1000$ Monte Carlo simulations with sample size $n=800$ per study, we compute the difference in ALTE across two studies and conduct a Wald test of equality between the two estimates. The results are reported in Table 1. The average cross-study gap is $0.256$ for PCA and $0.366$ for ICW. PCA rejects the true null hypothesis of equal ALTE in $24.7\%$ of replications, while ICW rejects it in $100\%$ of replications. As expected, because PCA and ICW are measurement-dependent methods, the resulting estimates are also measurement-dependent and therefore not comparable, even when the latent outcome and ALTE are identical across studies.
@@ 5 Application
Table 2: NSI and WSI estimates for the Kalla–Broockman application
@@
  • Notes: Entries are coefficient estimates with standard errors in parentheses. $^{*}p<0.10$, $^{**}p<0.05$, and $^{***}p<0.01$. The first three columns report NSI estimators using series, random-forest, and RKHS bases, respectively, to estimate the bridge component. The instrument set includes the two treatment indicators and baseline covariate. The last column reports the WSI estimator.
Table 2 reports the results. We use the two treatment assignment and baseline covariate as IVs in the first stage to estimate the bridge function. Panel A presents specifications in which the latent outcome is regressed on the two treatment indicators only. Panel B adds the baseline covariate to the second-stage regression. Across all three nonparametric implementations, the estimated effect of the full treatment is positive and statistically distinguishable from zero at conventional levels in most specifications. The point estimates are also quite stable across basis choices, ranging from $0.418$ to $0.444$ without covariate adjustment and from $0.398$ to $0.425$ with covariate adjustment. The abbreviated treatment has a smaller estimated effect and is not statistically distinguishable from zero in either panel. These findings mirror the substantive conclusion in Fu and Green (2025): the full canvassing intervention produces a meaningful shift in the latent immigration-attitude outcome, whereas the abbreviated treatment does not.
@@ 1. $r$ belongs to the orthogonal complement $\mathcal{N}(K^{*})^{\perp}$; and
We observe that $\mathbb{E}[\varphi_{j}(Y_{ij})|\eta_{i}]=\mathbb{E}[Y_{i1}|\eta_{i}]$ can be represented by $K\varphi=r$, where $r(\eta_{i}):=\mathbb{E}[Y_{i1}|\eta_{i}]\in H_{2}$ by regularity assumption i.
@@ B.2 Proof of Proposition 2
Now, let $K:\varphi\mapsto\mathbb{E}[\varphi\mid W_{i}]$ and define $r(W_{i})=\mathbb{E}[Y_{i1}\mid W_{i}]$. We obtain the well known NPIV equation $K\varphi=r$. The function $\varphi$ uniquely identified if and only if completeness condition hold: for any square-integrable function $g$, $\mathbb{E}[g(Y_{ij})|W_{i}]=0$ for almost all $W_{i}$ implies $g(Y_{ij})=0$ almost surely (Newey and Powell, 2003). ∎